# Differential geometry of particle motion in Stokesian regime

### v1

@@ I Particle trajectories and Geodesics
We start with the expression for energy dissipated along a trajectory of the particle (see Appendix A for a short derivation):
@@
$E=\int_{0}^{t}R_{ij}\frac{dx^{i}}{d\tau}\frac{dx^{j}}{d\tau}~d\tau.$
If one applies the Euler-Lagrange equations to extremize this dissipation functional, one obtains the standard homogeneous geodesic equations for the metric $g_{ij}=R_{ij}$. This suggests a compelling hypothesis: Are the physical trajectories of Stokesian particles geodesics of the resistance metric?
To test this, we examine the covariant acceleration of a particle driven by a constant external force $F_{k}$ (e.g., gravity). The equation of motion is given by the mobility relation $U^{i}=g^{ij}F_{j}$. Differentiating this with respect to time yields the kinematic acceleration:
@@
This result contradicts the hypothesis that the trajectory is a geodesic of the resistance metric alone. While the integral (2) describes the cost of motion, the physical particle driven by a constant Euclidean force does not traverse the path of minimum resistance. Instead, it experiences a geometric “drift” perpendicular to the geodesic path, necessitating a more unified geometric treatment.
@@ II The Jacobi-Maupertuis Resolution
The deviation observed in Equation (5) arises because the resistance metric $g_{ij}=R_{ij}$ describes the geometry of the medium, defining the instantaneous cost of motion, but ignores the dynamics of the driving force. To reconcile the kinematics with the geometry, we invoke a dissipative analogue to the Jacobi-Maupertuis principle.
@@
This resolves the discrepancy found in the pure resistance formulation. While the resistance tensor $R_{ij}$ creates a geometric “drift”, the scaling factor $\mathcal{D}(\mathbf{x})$ introduces a counter-acting “conformal force.” As derived explicitly in Appendix D, the gradient of the dissipation rate exactly cancels the geometric drift arising from the resistance tensor.
@@ III Application to a two-sphere system
$H(r)=\displaystyle\int_{r}^{\infty}\frac{R_{A}(s)-R_{B}(s)}{sR_{A}(s)}ds.$
@@ IV.2 The Duality of Stokesian Metrics
IV.3 Broader Implications
The concept of the Unified Dissipative Metric extends beyond simple sedimentation. Any low-Reynolds-number system where the local dissipation rate $\mathcal{D}$ can be defined as a function of position—including electrophoresis, magnetophoresis, or motion through porous media—can be mapped onto a Riemannian manifold. This opens the door to “Geometric Microfluidics,” where obstacle geometries are designed specifically to curve the effective spacetime of the channel, focusing or sorting particles based on their anisotropic resistance signatures.
@@ Acknowledgements.
Appendix A Power dissipated by a single moving particle in a quiescent fluid
$\dot{E}=\int_{V_{p}}\mathbf{\sigma}\cdot\mathbf{e}~dV,$ (12)
$E=\int_{0}^{t}\mathbf{R}\cdot\mathbf{U}\cdot\mathbf{U}~d\tau,$ (14)
Appendix B Geodesics on a manifold
$\ell_{P_{1},P_{2}}=\int_{s_{1}}^{s_{2}}~ds,$ (15)
where, dsds is the line element along the curve, defined as the quadratic form
$ds^{2}=g_{ij}dx^{i}dx^{j},$ (16)
if $g_{ij}$ is the metric tensor on a manifold $\mathcal{M}$.
= ∫_α_1^α_2 L(α)  dα
= ∫_α_1^α_2 g_ijdxi(α)dαdxj(α)dα dα.
Writing the Euler-Lagrange equations corresponding to equation (B) for Lagrangian $L(\alpha)$ give us the following (inhomogeneous) geodesic equation (Ch. 17, eq. 17.7 in [6]): {align} d2xidα2 + g^ij∂gjk∂xm dxkdα dxmdα - 12g^ij∂gkm∂xj dxkdα dxmdα
= 12dxidα(∂gkm∂xpdxkdα dxmdα dxpdα).
and consider the integral,
Note that the same curves on $\mathcal{M}$ minimize (19) as those which minimize (B), just with a different parametrization.
The parameters $\alpha$ and $\beta$ are related to each other as:
$\frac{d\beta}{d\alpha}=L(\alpha).$ (21)
Appendix C Proof of Geometric Consistency
@@
C.1 Transformation of the Covariant Acceleration
@@
C.2 The Dissipation Metric Proof
@@
C.3 Affine Parameterization and Energy Dissipation
@@
Since $\mathcal{D}$ represents the rate of energy dissipation, the differential dsds corresponds to the incremental energy dissipated along the path.
@@
C.4 Variational Derivation of the Geodesic Equation
@@
Appendix D Derivation of the Two-Sphere Solution
In this appendix, we explicitly re-derive the Risbud-Drazer trajectory, $b_{in}=y\exp{H(r)}$. Given that it is a system under Stokesian regime and the particle is moving under the action of a constant force, it trajectory is already proven to be a geodesic on the corresponding dissipation manifold.

### v2

@@ I Particle trajectories and Geodesics
We start with the expression for energy dissipated along a trajectory of the particle (see Appendix B for a short derivation):
@@
$E=\int_{0}^{t}R_{ij}\frac{dx^{i}}{d\tau}\frac{dx^{j}}{d\tau}~d\tau.$ (3)
Note that throughout this letter, we will use Einstein summation notation wherever applicable. Further, we will use the standard convention in differential geometry, wherein the lower indices of a tensor denote its covariant components, the upper indices denote the contravariant components, and the inverse of a covariant tensor $g_{ij}$ is the contravariant tensor $g^{ij}$.
If one applies the Euler-Lagrange equations to extremize the dissipation functional, one obtains the standard homogeneous geodesic equations for the metric $g_{ij}=R_{ij}$. This suggests a compelling hypothesis: Are the physical trajectories of Stokesian particles geodesics of the resistance metric?
To test this, we examine the covariant acceleration of a particle driven by a constant external force $F_{k}$ (e.g., gravity). The equation of motion is given by the mobility relation (in the standard differential geometric notation):
$U^{i}=(g_{ij})^{-1}F_{j}=g^{ij}F_{j}.$
Differentiating this with respect to time yields the kinematic acceleration:
@@
This result proves that the the particle trajectory is not a geodesic of the resistance metric alone.
[…]
Using a two-sphere system as an example (see §III), Figure 1 shows the geodesics obtained after integrating
$\frac{DU^{i}}{dt}=\frac{dU^{i}}{dt}+\Gamma^{i}_{jk}U^{j}U^{k}=0,$
such that $\Gamma^{i}_{jk}$ are the Christoffel symbols corresponding to $g_{ij}=R_{ij}$. Clearly, the geodesics do not show correct particle motion around a spherical obstacle of the same size. The metric does not have the correct curvature to “bend” the trajectories back to symmetry after crossing the obstacle. The reason is that the integral (2) describes only the cost of motion. A particle driven by a constant Euclidean force does not traverse the path of minimum resistance. Instead, it experiences a geometric “drift” from the geodesic path of the metric $g_{ij}=R_{ij}$, necessitating a more unified geometric treatment, which we describe next.
@@ II The Jacobi-Maupertuis Resolution
The unrealistic trajectories observed in Figure 1 (a result of Equation (6)) arise because the resistance metric $g_{ij}=R_{ij}$ describes the geometry of the medium, defining the instantaneous cost of motion, but ignores the dynamics of the driving force. To reconcile the kinematics with the geometry, we invoke a dissipative analogue to the Jacobi-Maupertuis principle.
@@
This resolves the discrepancy found in the pure resistance formulation. While the resistance tensor $R_{ij}$ creates a geometric “drift”, the scaling factor $\mathcal{D}(\mathbf{x})$ introduces a counter-acting “conformal force.” As derived explicitly in Appendix A, the gradient of the dissipation rate exactly cancels the geometric drift arising from the resistance tensor. The geodesics in Figure 2 show this effect for the two-sphere system, as contrasted with Figure 1.
@@
Prima facie, it appears to be paradoxical that memoryless Stokesian dynamics governed by a local kinematic relation $\mathbf{U}=\mathbf{M}(\mathbf{x})\cdot\mathbf{F}$ also satisfies a global variational problem with the cumulative dissipated energy, $s=\int\mathcal{D}dt$ as the affine parameter. In other words, how can a particle with no memory navigate a path of globally stationary dissipative action?
This is analogous to Fermat’s Principle of Least Time in optics: a photon propagating through a continuously varying dielectric medium has no memory. At every infinitesimal step, its propagation is governed strictly by the local properties of the medium. The continuous sequence of the strictly local, memoryless steps naturally traces out a trajectory that globally minimizes the travel time.
Similarly, in our geometric formulation, the particle locally moves in the direction dictated by the local mobility tensor $\mathbf{M}(\mathbf{x})$ at every instant. However, the mobility tensor itself is ‘shaped’ by the global boundary conditions and the geometry of the problem. This leads to a curved “dissipation landscape.” The mathematical equivalence between the first-order local kinematic equation and the second-order geodesic equation on $\tilde{\mathbf{g}}$ demonstrates that obeying the local resistance gradient is identical to tracing a geodesic in this curved space.
[…]
@@ III Application to a two-sphere system
$H(r)=\displaystyle\int_{r}^{\infty}\frac{R_{B}(s)-R_{A}(s)}{sR_{B}(s)}ds.$
@@
Figure 2 shows a perfect match between Risbud-Drazer trajectories and the geodesics of $\mathbf{\tilde{g}}$ for the two-sphere problem, thus validating the differential geometric framework for this system.
@@ IV.2 The Duality of Stokesian Metrics
IV.3 Broader Implications and Future Directions
The concept of the Unified Dissipative Metric extends well beyond the simple two-sphere scattering problem, offering both conceptual shifts as well as practical computational advantages.
First, as the problem geomtry gets more complicated, framing the kinematics as a Riemannian manifold problem enables us to use the extensive analytical and computational machinery of differential geometry. In complex microfluidic environments—such as porous media or anisotropic polymer networks—the spatial resistance tensor $\mathbf{R}(\mathbf{x})$ can be mapped numerically via standard Stokes solvers. Once the metric field $\tilde{\mathbf{g}}=\mathcal{D}\mathbf{R}$ is established, one can leverage standard, open-source differential geometry packages (such as geomstats or EinsteinPy in Python, or the DifferentialGeometry suite in Maple) to directly compute the geodesics. This translates trajectory prediction from an iterative, time-stepping simulation into a robust geometric boundary-value problem. Furthermore, this perspective enables “Geometric Microfluidics,” where the spatial layout of fixed obstacles can be deliberately designed to sculpt the manifold’s curvature, creating gravitational-like “lenses” to focus, sort, or trap particles based on their specific anisotropic resistance signatures.
Second, regarding many-body systems: the present work restricts itself to a single particle navigating a fixed obstacle field, defining a 3D manifold. For a suspension of $N$ hydrodynamically interacting particles, the grand resistance tensor naturally defines the metric on a $6N$-dimensional configuration manifold. Conceptually, the collective evolution of the entire suspension is described by a single geodesic in this hyper-space. While analytically computing Christoffel symbols in $6N$ dimensions is computationally intractable for large $N$, this global geometric viewpoint provides a new theoretical lens for statistical mechanics. Just as the geometric curvature of phase space dictates ergodicity in Hamiltonian systems, the scalar curvature of this high-dimensional dissipative manifold could yield novel, macro-scale bounds on the bulk rheology and collective transport properties of concentrated suspensions.
Finally, this geometric formalism establishes a rigorous foundation for exploring stochastic transport. This Letter explicitly limits its scope to the deterministic, non-Brownian regime. However, in the colloidal domain where thermal noise is dominant, deterministic paths are replaced by stochastic trajectory distributions. A highly promising avenue for future research is to couple the Unified Dissipative Metric with the covariant formulation of the Onsager-Machlup functional. We are currently investigating whether the “most probable trajectory” of a Brownian Stokesian particle coincides exactly with the deterministic geodesic of $\tilde{\mathbf{g}}$, or if the intrinsic curvature of the resistance manifold induces a purely geometric, noise-driven drift.
@@ Acknowledgements.
Appendix A Two Proofs that trajectories are geodesics of the metric $\mathbf{\tilde{g}}=\mathcal{D}\mathbf{R}$
@@
A.1 A Proof Using Conformal Scaling and Affine Parametrization
A.1.1 Transformation of the Covariant Acceleration
@@
A.1.2 The Dissipation Metric Proof
@@
A.1.3 Affine Parameterization and Energy Dissipation
@@
Since $\mathcal{D}$ represents the rate of energy dissipation, the differential $ds$ corresponds to the incremental energy dissipated along the path.
@@
A.2 A Proof Using the Variational Derivation of the Geodesic Equation
@@
Appendix B A Short Derivation of Equation (2): Single Particle Dissipation
[Note: This derivation is well-known and presented here for the sake of completeness.]
$\dot{E}=\int_{V_{p}}\mathbf{\sigma}\cdot\mathbf{e}~dV,$ (32)
$E=\int_{0}^{t}\mathbf{R}\cdot\mathbf{U}\cdot\mathbf{U}~d\tau,$ (34)
Appendix C A Brief Introduction to Geodesics on a manifold
[Note: This appendix is presented for the convenience of a reader uninitiated in differential geometry. The results are well-known and available in any textbook of differential geometry or general relativity.]
$\ell_{P_{1},P_{2}}=\int_{s_{1}}^{s_{2}}~ds,$ (35)
where, $ds$ is the line element along the curve, defined as the quadratic form
$ds^{2}=g_{ij}dx^{i}dx^{j},$ (36)
if $g_{ij}$ is the metric tensor on a manifold $\mathcal{M}$.
= ∫_α_1^α_2 L(α)  dα
= ∫_α_1^α_2 g_ijdxi(α)dαdxj(α)dα dα.
Writing the Euler-Lagrange equations corresponding to equation (C) for Lagrangian $L(\alpha)$ give us the following (inhomogeneous) geodesic equation (Ch. 17, eq. 17.7 in [6]): {align} d2xidα2 + g^ij∂gjk∂xm dxkdα dxmdα - 12g^ij∂gkm∂xj dxkdα dxmdα
= 12dxidα(∂gkm∂xpdxkdα dxmdα dxpdα).
and consider the integral,
Note that the same curves on $\mathcal{M}$ minimize (39) as those which minimize (C), just with a different parametrization.
The parameters $\alpha$ and $\beta$ are related to each other as:
$\frac{d\beta}{d\alpha}=L(\alpha).$ (41)
Appendix D Re-Derivation of the Two-Sphere Solution
For the convenience of a reader, here we explicitly re-derive the Risbud-Drazer trajectory, $b_{in}=y\exp{H(r)}$. Given that it is a system under Stokesian regime and the particle is moving under the action of a constant force, it trajectory is already proven to be a geodesic on the corresponding dissipation manifold. A comparison between the equation of the trajectory derived here and the solution to the geodesic equation is shown in Figure 2.